# End-to-End Breast Cancer Radiotherapy Planning via LMMs with Consistency Embedding

### v3

@@ IV-B Implementation details
LLaMA2 [23] - 0.625 (0.603-0.646) 0.439 (0.423-0.454) 0.625 (0.603-0.646) 0.667 (0.649-0.684) 0.468 (0.456-0.479) 0.667 (0.649-0.684)
MedAlpaca [25] - 0.529 (0.484-0.578) 0.367 (0.334-0.399) 0.529 (0.484-0.578) 0.414 (0.362-0.462) 0.329 (0.283-0.373) 0.414 (0.362-0.462)
ChatDoctor [24] - 0.617 (0.572-0.657) 0.411 (0.381-0.439) 0.617 (0.572-0.657) 0.549 (0.508-0.588) 0.425 (0.392-0.459) 0.549 (0.508-0.588)
Asclepius [27] - 0.662 (0.640-0.682) 0.447 (0.436-0.458) 0.662 (0.640-0.682) 0.590 (0.571-0.609) 0.412 (0.401-0.423) 0.590 (0.571-0.609)
PMC-LLaMA [26] - 0.398 (0.341-0.451) 0.239 (0.200-0.276) 0.398 (0.341-0.451) 0.382 (0.327-0.434) 0.235 (0.197-0.275) 0.382 (0.327-0.434)
ChatGPT - 0.672 (0.641-0.703) 0.467 (0.447-0.486) 0.672 (0.641-0.703) 0.651 (0.615-0.683) 0.456 (0.434-0.476) 0.651 (0.615-0.683)
RO-LMM-S Vanila 0.798 (0.777-0.818) 0.616 (0.598-0.635) 0.798 (0.777-0.818) 0.788 (0.769-0.806) 0.607 (0.586-0.628) 0.788 (0.769-0.806)
Therapy Plan Suggestion
LLaMA2 [23] - 0.077 (0.071-0.083) 0.038 (0.034-0.042) 0.077 (0.071-0.083) 0.069 (0.065-0.073) 0.029 (0.027-0.031) 0.069 (0.065-0.073)
MedAlpaca [25] - 0.139 (0.126-0.152) 0.053 (0.049-0.058) 0.139 (0.126-0.152) 0.137 (0.123-0.152) 0.048 (0.044-0.053) 0.137 (0.123-0.152)
ChatDoctor [24] - 0.214 (0.194-0.234) 0.082 (0.072-0.091) 0.214 (0.194-0.234) 0.191 (0.178-0.205) 0.065 (0.060-0.071) 0.191 (0.178-0.205)
Asclepius [27] - 0.176 (0.162-0.190) 0.075 (0.067-0.083) 0.176 (0.162-0.190) 0.142 (0.134-0.149) 0.055 (0.051-0.059) 0.142 (0.134-0.149)
PMC-LLaMA [26] - 0.118 (0.101-0.137) 0.034 (0.029-0.040) 0.118 (0.101-0.137) 0.165 (0.136-0.195) 0.050 (0.040-0.060) 0.165 (0.136-0.195)
GPT4.0 - 0.366 (0.343-0.391) 0.230 (0.207-0.252) 0.366 (0.343-0.391) 0.316 (0.294-0.339) 0.249 (0.226-0.272) 0.316 (0.294-0.339)
RO-LMM-P Vanila 0.651 (0.617-0.689) 0.486 (0.437-0.539) 0.651 (0.617-0.689) 0.591 (0.574-0.609) 0.426 (0.403-0.452) 0.591 (0.574-0.609)
RO-LMM-P+ NEFT 0.639 (0.607-0.673) 0.467 (0.419-0.516) 0.639 (0.607-0.673) 0.588 (0.573-0.606) 0.419 (0.399-0.441) 0.588 (0.573-0.606)
RO-LMM-P++ CEFT 0.668 (0.632-0.705) 0.512 (0.460-0.564) 0.668 (0.632-0.705) 0.615 (0.592-0.638) 0.459 (0.429-0.489) 0.615 (0.592-0.638)
TABLE III: Clinical expert analysis for report summarization. R#: each rubric, C#:each clinical expert.
@@
C2 6.0 / 3.0 / 9.0 / 7.7 / 25.0 8.0 / 0.0 / 0.0 / 9.0 / 17.0 10.0 / 9.0 / 8.0 / 9.0 / 36.0
IV-C Evaluation Metric and Baseline Models
The evaluation of generated clinical report summaries and treatment plans utilizes common NLP metrics, including Rouge-1, Rouge-2, and Rouge-L [47]. However, acknowledging potential limitations in domain specificity for both tasks, we conduct evaluations using two clinical experts and GPT-4.0. A board-certified expert creates five scoring rubrics for plan suggestions, detailed in Table A1. Using these rubrics, the experts score each aspect of the plans, assigning 0 or 1 for each rubric separately. Similarly, for clinical report summarization, rubrics are created and the same evaluation process is followed. Inspired by the recent evaluation methodologies such as GPT-Score [48] and G-Eval [49], GPT-4.0 also evaluates specific examples of outputs based on these rubrics with in-context learning on the treatment plan task. To evaluate the 3D target volume segmentation performance using a ground truth segmentation mask, we measure Dice coefficient (Dice), Intersection over Union (IoU), and the 95th percentile of Hausdorff Distance (HD-95) [50] in centimeter (cm) unit to measure spatial distances between the ground truth masks and the predicted segmentation results.
For text-related tasks, to validate the effectiveness of our model, we compare our model with several clinical LLMs serving as baselines, specifically MedAlpaca [25], ChatDoctor [24], Asclepius [27], and PMC-LLAMA [26]. Additionally, we include a comparison with ChatGPT [16] for the report summarization task and GPT-4.0 for the treatment plan suggestion task using few-shot in-context learning.
For the segmentation task, we utilize 3D Residual U-Net [42] and ConTEXTualNET [37] as our baseline methods. We further utilize an additional baseline as a variant of RO-LMM-SEG, namely LLaMA2-SEG, which is initialized with the checkpoint of vanilla LLaMA-2-7B-Chat model.
TABLE IV: Clinical expert analysis for treatment plan suggestion. R#: each rubric, C#: each clinical expert.
@@ IV-C Evaluation Metric and Baseline Models
TABLE V: Comparison of 3D target volume segmentation performance.
@@
Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI) Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI)
3D Residual U-Net [42] - 0.804 (0.777-0.831) 0.688 (0.653-0.720) 9.018 (7.077-11.170) 0.690 (0.652-0.725) 0.550 (0.511-0.588) 25.057 (22.844-27.070)
$\text{ConTEXTualNET}^{\dagger}$ [37] - 0.812 (0.783-0.838) 0.698 (0.661-0.731) 7.392 (5.793-9.083) 0.696 (0.663-0.726) 0.551 (0.517-0.583) 21.253 (20.002-22.474)
LLaMA2-SEG - 0.817 (0.790-0.841) 0.705 (0.670-0.734) 9.101 (7.189-11.129) 0.692 (0.666-0.714) 0.539 (0.514-0.560) 24.087 (22.450-25.789)
RO-LMM-SEG Vanilla 0.824 (0.799-0.846) 0.713 (0.680-0.742) 5.528 (3.978-7.295) 0.778 (0.743-0.806) 0.655 (0.621-0.686) 17.459 (15.254-19.678)
RO-LMM-SEG+ NESEG 0.829 (0.805-0.851) 0.719 (0.690-0.749) 4.636 (3.351-6.073) 0.781 (0.741-0.815) 0.664 (0.626-0.699) 29.410 (27.230-31.386)
RO-LMM-SEG++ CESEG 0.840 (0.818-0.860) 0.734 (0.706-0.760) 4.391 (3.227-5.915) 0.766 (0.728-0.797) 0.642 (0.607-0.673) 15.236 (12.807-17.536)
modified for 3D target volume segmentation with LLaMA2-7B-Chat as backbone for fair comparison
TABLE VI: Comparison of 3D target segmentation performance for overall and specific patient types.
@@
#Patients Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI) #Patients Dice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI)
All Unimodal 79 0.804 (0.777-0.831) 0.688 (0.653-0.720) 9.018 (7.077-11.170) 81 0.690 (0.652-0.725) 0.550 (0.511-0.588) 25.057 (22.844-27.070)
Multimodal 0.840 (0.818-0.860) 0.734 (0.706-0.760) 4.391 (3.227-5.915) 0.766 (0.728-0.797) 0.642 (0.607-0.673) 15.236 (12.807-17.536)
Gain $\uparrow$ +0.036 +0.046 -4.627 +0.076 +0.092 -9.821
Typical Unimodal 55 0.842 (0.821-0.861) 0.734 (0.705-0.761) 7.374 (5.447-9.687) 73 0.723 (0.694-0.749) 0.579 (0.546-0.610) 25.082 (22.783-27.380)
Multimodal 0.877 (0.866-0.888) 0.784 (0.766-0.802) 3.200 (2.361-4.303) 0.782 (0.742-0.813) 0.659 (0.621-0.690) 14.716 (12.407-17.204)
Gain $\uparrow$ +0.035 +0.049 -4.173 +0.059 +0.080 -10.366
Atypical Unimodal 24 0.720 (0.654-0.784) 0.584 (0.513-0.657) 12.519 (8.553-16.584) 8 0.370 (0.178-0.594) 0.259 (0.106-0.449) 27.138 (22.417-30.807)
Multimodal 0.760 (0.701-0.807) 0.628 (0.562-0.682) 6.948 (3.820-11.014) 0.585 (0.346-0.753) 0.458 (0.269-0.610) 22.410 (16.377-28.314)
Gain $\uparrow$ +0.040 +0.044 -5.571 +0.215 +0.199 -4.728
IV-D Results
Clinical Report Summarization. We present the performance of our model on the clinical report summarization task, along with confidence intervals for each method, in Table II. Our fine-tuned model of RO-LMM-S demonstrate significant improvements over the baselines, providing consistent margins in all metrics and confidence intervals. Notably, RO-LMM-S outperforms ChatGPT with few-shot in-context learning.
Moreover, we evaluate the generated summaries using expertise-based rubrics by two clinical experts and compare them to baselines, including ChatGPT and LLaMa-2. As shown in Table III, our ROB-LMM-S model significantly outperforms all baselines in both internal and external validations, thanks to its domain-specific knowledge. Additionally, Pearson correlation ($r$) analysis reveals strong positive inter-clinician correlations ( $>0.85$ and $>0.95$ for internal and external validation, respectively), confirming the reliability of our rubrics and the clinical relevance of ROB-LMM-S. Therefore, our ROB-LMM-S provides practical and meaningful summaries that can assist in the field of radiation oncology.
Treatment Plan Suggestion. As indicated in Table II, we evaluate the model’s performance on treatment plan suggestion task based on generated clinical notes. Similar to the report summarization task, our fine-tuned variants, RO-LMM-P, exhibit substantial enhancements in suggestion performance compared to other baselines. Notably, RO-LMM-P+ shows inferior performance on the external dataset compared to the vanilla approach. However, our proposed RO-LMM-P++ achieves the best performance on both internal and external datasets across all metrics.
Furthermore, we comprehensively evaluate the generated plans using our expertise-based rubrics and clinical expert analysis, as detailed in Table IV. For both internal and external validation, our RO-LMM variants significantly outperform GPT-4.0 with few-shot in-context learning, as assessed by both GPT and clinical experts, due to their domain-specific knowledge. We confirm that our model with CEFT achieved the best performance in terms of clinical evaluation, especially on the external set, demonstrating clinical effectiveness in line with our goals. We further evaluate inter-clinicians correlation and the correlation between clinicians and GPT-4, using Pearson correlation ($r$). The results show strong positive correlations ($>0.7$ and $>0.4$, for each), affirming reliability of our proposed evaluation rubrics.
3D Target Volume Segmentation. As reported in Table V, our proposed plan-guided 3D target segmentation frameworks excel other baseline methods with large gains, specifically remained stable performance across the external validation setting. This can be also observed in qualitative results. As shown in Figure 3(a) and (c), despite the ground truth mask posing target volume on the one side of the breast, the unimodal 3D Residual U-Net incorrectly segments contours not only the target volume but also the outside of the target breast, while all the multimodal RO-LMM variants correctly segment target volume aligning consistently with the ground truth masks. Moreover, as shown in Figure 3(a) and (b), RO-LMM-SEG+ with NESEG module yields noisy segmentation output as indicated as red arrows, and RO-LMM-SEG sometimes under-segments the target volume. In contrast, our proposed RO-LMM-SEG++ with CESEG module accurately contours the breast and regional lymph nodes that need to be treated. We further categorize all patients into two distinct types: patients those who had breast-conserving surgery (typical type) and patients those who underwent total mastectomy (atypical type). The comparative results are presented in the Table VI separately, with the unimodal model referred to the 3D Residual U-Net and the multimodal model denoted as our proposed RO-LMM-SEG++ with CESEG. Firstly, our proposed multimodal model shows substantial performance improvements of approximately up to 5% and 10% for internal and external validation, respectively. When comparing the typical and atypical types of patients, particularly in the less prevalent atypical type during external validation, we observed notable enhancements with a performance gain of up to 20%. This substantial generalizbility observed in extremely low-prevalence distributions demonstrates the data efficiency of our multimodal concept, leveraging textual information compared to the traditional unimodal approach.
Inference on Public Dataset. To validate the effectiveness of our methods when AI-generated results are reused as input, we evaluated the performance by inputting each model’s report summarization results its plan suggestion task. We conducted experiments on a publicly available text-related dataset, as detailed in Table VII. Our proposed methods consistently outperformed other specialized medical LLMs and variants of GPT (such as ChatGPT and GPT-4) in both clinical report summarization and therapy plan suggestion tasks. Notably, our RO-LMM-P++ method demonstrated substantial performance enhancements, with improvements of up to 0.55 and 0.31 over other medical LLMs and GPT variants, respectively, demonstrating our model’s robustness to error accumulation during sequential generations.
@@ IV-D Results
TABLE VII: Quantitative comparison results for our RO-LMM’s clinical report summarization and therapy plan suggestion performance on the publicly available dataset.
@@
Rouge 1 $\uparrow$ (CI) Rouge 2 $\uparrow$ (CI) Rouge L $\uparrow$ (CI)
@@
Therapy Plan Suggestion

### v4

@@ 4.1 Dataset
LLaMA2 [46] - 0.621 (0.601-0.642) 0.434 (0.419-0.449) 0.621 (0.601-0.642) 0.667 (0.649-0.684) 0.468 (0.456-0.479) 0.667 (0.649-0.684)
MedAlpaca [15] - 0.514 (0.460-0.549) 0.351 (0.318-0.392) 0.514 (0.460-0.549) 0.414 (0.362-0.462) 0.329 (0.283-0.373) 0.414 (0.362-0.462)
ChatDoctor [29] - 0.601 (0.544-0.641) 0.400 (0.370-0.424) 0.601 (0.544-0.641) 0.549 (0.508-0.588) 0.425 (0.392-0.459) 0.549 (0.508-0.588)
Asclepius [25] - 0.652 (0.624-0.674) 0.442 (0.429-0.452) 0.652 (0.624-0.674) 0.590 (0.571-0.609) 0.412 (0.401-0.423) 0.590 (0.571-0.609)
PMC-LLaMA [50] - 0.390 (0.334-0.436) 0.233 (0.194-0.270) 0.390 (0.334-0.436) 0.382 (0.327-0.434) 0.235 (0.197-0.275) 0.382 (0.327-0.434)
ChatGPT - 0.657 (0.614-0.685) 0.456 (0.434-0.473) 0.657 (0.614-0.685) 0.651 (0.615-0.683) 0.456 (0.434-0.476) 0.651 (0.615-0.683)
RO-LMM-S Default 0.783 (0.758-0.806) 0.605 (0.582-0.625) 0.783 (0.758-0.806) 0.788 (0.769-0.806) 0.607 (0.586-0.628) 0.788 (0.769-0.806)
Radiotherapy Strategy Suggestion
LLaMA2 [46] - 0.075 (0.068-0.082) 0.036 (0.032-0.040) 0.075 (0.068-0.082) 0.069 (0.065-0.073) 0.029 (0.027-0.031) 0.069 (0.065-0.073)
MedAlpaca [15] - 0.133 (0.120-0.145) 0.051 (0.047-0.055) 0.133 (0.120-0.145) 0.137 (0.123-0.152) 0.048 (0.044-0.053) 0.137 (0.123-0.152)
ChatDoctor [29] - 0.205 (0.184-0.225) 0.079 (0.069-0.087) 0.205 (0.184-0.225) 0.191 (0.178-0.205) 0.065 (0.060-0.071) 0.191 (0.178-0.205)
Asclepius [25] - 0.169 (0.155-0.181) 0.072 (0.064-0.079) 0.169 (0.155-0.181) 0.142 (0.134-0.149) 0.055 (0.051-0.059) 0.142 (0.134-0.149)
PMC-LLaMA [50] - 0.113 (0.096-0.131) 0.033 (0.028-0.038) 0.113 (0.096-0.131) 0.165 (0.136-0.195) 0.050 (0.040-0.060) 0.165 (0.136-0.195)
GPT4.0 - 0.356 (0.326-0.375) 0.222 (0.197-0.240) 0.356 (0.326-0.375) 0.316 (0.294-0.339) 0.249 (0.226-0.272) 0.316 (0.294-0.339)
RO-LMM-P Default 0.634 (0.590-0.670) 0.471 (0.418-0.521) 0.634 (0.590-0.670) 0.591 (0.574-0.609) 0.426 (0.403-0.452) 0.591 (0.574-0.609)
RO-LMM-P+ NEFT 0.623 (0.576-0.660) 0.453 (0.399-0.510) 0.623 (0.576-0.660) 0.588 (0.573-0.606) 0.419 (0.399-0.441) 0.588 (0.573-0.606)
RO-LMM-P++ CEFT 0.655 (0.616-0.693) 0.500 (0.449-0.554) 0.655 (0.616-0.693) 0.615 (0.592-0.638) 0.459 (0.429-0.489) 0.615 (0.592-0.638)
Table 3: Clinical expert analysis for report summarization. R#: each rubric, C#:each clinical expert.
@@
C2 6.0 / 3.0 / 9.0 / 7.0 / 25.0 8.0 / 0.0 / 0.0 / 9.0 / 17.0 10.0 / 9.0 / 8.0 / 9.0 / 36.0
4.2 Implementation details
For instruction fine-tuning for RO-LMM-S and RO-LMM-P, we use the LLaMA-2-7B-Chat [46] model as a Default. We adopt full-finetuning approach for all of our models. The maximum context length is set at 4096, and the batch size is set to 2 for each GPU, leveraging 4 NVIDIA A6000 GPUs. We use the AdamW [35] optimizer, with a learning rate of 5e-5 until reaching 3 epochs for both RO-LMM-S and RO-LMM-P. The hyper-parameter $\lambda$ for CEFTune is set to 1 in all of tasks, and for consistency regularization we adopt the variants of SBERT [42] which is trained on PubMed [^2] https://www.ncbi.nlm.nih.gov/pubmed/ dataset, called PubMedBERT.333www.huggingface.co/NeuML/pubmedbert-base-embeddings. To ensure fair comparisons with other approaches, we utilize the Fully Sharded Data Parallel (FSDP) acceleration method, the bf16 (Brain Floating Point) data format, and gradient checkpointing [4] in all experiments.
For training the segmentation network, we adopt the 3D Residual U-Net [8] as the backbone architecture, implemented using the MONAI open-source library [^4] https://monai.io/. During data preprocessing, all chest CT scans and CTV annotations are re-sampled to a uniform voxel spacing of 1.0 $\times$ 1.0 $\times$ 3.0 mm3. Image intensities are truncated to a Hounsfield Unit (HU) range of -1,000 to 1,000 and normalized linearly to fall within a scale of 0 to 1.0. For network training, 3D patches sized 384 $\times$ 384 $\times$ 128 pixels are randomly cropped to encompass the entire breast, using a batch size of 1 per GPU. Training utilizes 4 NVIDIA A6000 GPUs. Performance of the backbone is benchmarked against state-of-the-art models such as 3D SegMsamba [52] and 3D UNETR [16]. Due to memory constraints, input sizes of 384 $\times$ 384 $\times$ 128 pixels and 320 $\times$ 320 $\times$ 64 pixels are used for SegMamba and UNETR, respectively. For evaluation, the complete 3D CT volumes are processed using a sliding window approach across all models. To incorporate textual information into the RO-LMM-SEG/SEG+/SEG++ frameworks, we initialize the LLaMA-2-7B-Chat model with the pre-trained RO-LMM-P++ checkpoint. For training all the segmentation models, optimization is performed with the AdamW optimizer [35], using a learning rate of 1e-4 over 100 epochs.
Table 4: Clinical expert analysis for radiotherapy strategy suggestion. R#: each rubric, C#: each clinical expert.
@@ 4.2 Implementation details
4.3 Evaluation Metric and Default Models
The evaluation of generated clinical report summaries and radiotherapy strategies utilizes common NLP metrics, including Rouge-1 (R-1), Rouge-2 (R-2), and Rouge-L (R-L) [32].
However, acknowledging potential limitations in domain specificity for both tasks, we conduct evaluations using two clinical experts and GPT-4.0. A board-certified expert creates five scoring rubrics for plan suggestions, detailed in Table A2. Using these rubrics, the experts score each aspect of the plans, assigning 0 or 1 for each rubric separately. Similarly, for clinical report summarization, rubrics are created and the same evaluation process is followed. Inspired by the recent evaluation methodologies such as GPT-Score [14] and G-Eval [34], GPT-4.0 also evaluates specific examples of outputs based on these rubrics with in-context learning on the radiotherapy strategy suggestion task. To evaluate the 3D target volume segmentation performance using a ground truth segmentation mask, we measure Dice coefficient (Dice), Intersection over Union (IoU), and the 95th percentile of Hausdorff Distance (HD-95) [10] in millimeter (mm) unit to measure spatial distances between the ground truth masks and the predicted segmentation results.
For text-related tasks, to validate the effectiveness of our model, we compare our model with several clinical LLMs serving as Defaults, specifically MedAlpaca [15], ChatDoctor [29], Asclepius [25], and PMC-LLAMA [50]. Additionally, we include a comparison with ChatGPT [38] for the report summarization task and GPT-4.0 for the radiotherapy strategy suggestion task using few-shot in-context learning.
For the segmentation task, we utilize 3D Residual U-Net [8] and ConTEXTualNET [19] as our Default methods. We further utilize an additional Default as a variant of RO-LMM-SEG, namely LLaMA2-SEG, which is initialized with the checkpoint of vanilla LLaMA-2-7B-Chat model.
5 Results
5.1 Clinical Report Summarization.
We present the performance of our model on the clinical report summarization task, along with confidence intervals for each method, in Table 2. Our fine-tuned model of RO-LMM-S demonstrate significant improvements over the Defaults, providing consistent margins in all metrics and confidence intervals. Notably, RO-LMM-S outperforms ChatGPT with few-shot in-context learning.
Moreover, we evaluate the generated summaries using expertise-based rubrics by two clinical experts and compare them to Defaults, including ChatGPT and LLaMa-2. As shown in Table 3, our RO-LMM-S model significantly outperforms all Defaults in both internal and external validations, thanks to its domain-specific knowledge. Additionally, Pearson correlation ($r$) analysis reveals strong positive inter-clinician correlations ( $>0.85$ and $>0.95$ for internal and external validation, respectively), confirming the reliability of our rubrics and the clinical relevance of RO-LMM-S. Therefore, our RO-LMM-S provides practical and meaningful summaries that can assist in the field of radiation oncology.
5.2 Radiotherapy Strategy Suggestion.
As indicated in Table 2, we evaluate the model’s performance on radiotherapy strategy suggestion task based on generated clinical notes. Similar to the report summarization task, our fine-tuned variants, RO-LMM-P, exhibit substantial enhancements in suggestion performance compared to other Defaults. Notably, RO-LMM-P+ shows inferior performance on the external dataset compared to the vanilla approach. However, our proposed RO-LMM-P++ achieves the best performance on both internal and external datasets across all metrics.
Furthermore, we comprehensively evaluate the generated plans using our expertise-based rubrics and clinical expert analysis, as detailed in Table 4. For both internal and external validation, our RO-LMM variants significantly outperform GPT-4.0 with few-shot in-context learning, as assessed by both GPT and clinical experts, due to their domain-specific knowledge. We confirm that our model with CEFT achieved the best performance in terms of clinical evaluation, especially on the external set, demonstrating clinical effectiveness in line with our goals. We further evaluate inter-clinicians correlation and the correlation between clinicians and GPT-4, using Pearson correlation ($r$). The results show strong positive correlations ($>0.7$ and $>0.4$, for each), affirming reliability of our proposed evaluation rubrics.
Fig. 3: Qualitative comparison on 3D target volume segmentation task. Red arrows indicate errors.
Table 5: Comparison of 3D target volume segmentation performance.
@@ 5 Results
Dice $\uparrow$ (CI) surfaceDice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI) Dice $\uparrow$ (CI) surfaceDice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI)
3D Residual U-Net [8] - 0.802 (0.775-0.826) 0.517 (0.485-0.545) 0.684 (0.650-0.714) 45.852 (32.441-61.029) 0.689 (0.652-0.723) 0.409 (0.381-0.436) 0.548 (0.510-0.584) 132.825 (113.296-151.525)
3D SegMamba [52] - 0.576 (0.554-0.598) 0.364 (0.344-0.386) 0.411 (0.389-0.434) 146.610 (140.391-152.103) 0.569 (0.540-0.595) 0.315 (0.295-0.334) 0.408 (0.381-0.434) 156.767 (152.551-160.849)
3D UNETR [16] - 0.564 (0.544-0.584) 0.340 (0.319-0.361) 0.398 (0.379-0.419) 147.310 (141.625-152.392) 0.538 (0.523-0.551) 0.302 (0.287-0.316) 0.370 (0.357-0.382) 158.272 (154.231-162.241)
LLMSeg [37] - 0.814 (0.790-0.836) 0.531 (0.499-0.561) 0.699 (0.668-0.727) 42.814 (29.983-56.439) 0.681 (0.656-0.700) 0.397 (0.373-0.420) 0.525 (0.502-0.544) 160.151 (150.742-169.029)
$\text{ConTEXTualNET}^{\dagger}$ [19] - 0.774 (0.728-0.811) 0.491 (0.451-0.525) 0.664 (0.618-0.703) 31.772 (21.484-44.095) 0.745 (0.708-0.776) 0.444 (0.413-0.472) 0.622 (0.586-0.654) 82.013 (64.103-100.648)
RO-LMM-SEG Default 0.818 (0.795-0.841) 0.524 (0.493-0.555) 0.705 (0.676-0.736) 20.611 (13.502-29.389) 0.747 (0.715-0.775) 0.419 (0.388-0.449) 0.615 (0.580-0.646) 100.562 (86.656-115.053)
RO-LMM-SEG+ NESEG 0.828 (0.806-0.850) 0.518 (0.489-0.547) 0.718 (0.690-0.745) 17.805 (10.818-26.427) 0.774 (0.733-0.811) 0.458 (0.427-0.486) 0.659 (0.619-0.695) 34.487 (23.313-46.962)
RO-LMM-SEG++ CESEG 0.838 (0.818-0.858) 0.555 (0.523-0.583) 0.732 (0.706-0.757) 13.658 (8.133-20.770) 0.761 (0.723-0.794) 0.418 (0.391-0.443) 0.636 (0.601-0.668) 29.459 (21.433-38.557)
modified for 3D target volume segmentation
Table 6: Comparison of 3D target segmentation performance for overall and specific patient types.
@@
#Patients Dice $\uparrow$ (CI) surfaceDice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI) #Patients Dice $\uparrow$ (CI) surfaceDice $\uparrow$ (CI) IoU $\uparrow$ (CI) HD-95 $\downarrow$ (CI)
All Unimodal 79 0.802 (0.775-0.826) 0.517 (0.485-0.545) 0.684 (0.650-0.714) 45.852 (32.441-61.029) 81 0.689 (0.652-0.723) 0.409 (0.381-0.436) 0.548 (0.510-0.584) 132.825 (113.296-151.525)
Multimodal 0.838 (0.818-0.858) 0.555 (0.523-0.583) 0.732 (0.706-0.757) 13.658 (8.133-20.770) 0.761 (0.723-0.794) 0.418 (0.391-0.443) 0.636 (0.601-0.668) 29.459 (21.433-38.557)
Gain $\uparrow$ +0.036 +0.038 +0.048 -29.194 +0.072 +0.009 +0.088 -103.366
Typical Unimodal 55 0.837 (0.812-0.860) 0.553 (0.521-0.583) 0.728 (0.694-0.760) 37.449 (21.053-55.604) 73 0.720 (0.691-0.745) 0.429 (0.403-0.454) 0.576 (0.542-0.606) 133.108 (113.319-155.072)
Multimodal 0.876 (0.865-0.888) 0.603 (0.576-0.630) 0.783 (0.765-0.799) 7.385 (6.095-8.708) 0.776 (0.737-0.808) 0.427 (0.398-0.456) 0.652 (0.615-0.684) 27.519 (19.673-36.031)
Gain $\uparrow$ +0.039 +0.050 +0.055 -30.064 +0.056 -0.002 +0.076 -105.589
Atypical Unimodal 24 0.722 (0.658-0.781) 0.435 (0.374-0.494) 0.585 (0.516-0.652) 67.873 (36.956-98.711) 8 0.403 (0.218-0.593) 0.225 (0.108-0.349) 0.292 (0.132-0.460) 128.913 (91.743-157.660)
Multimodal 0.758 (0.702-0.801) 0.447 (0.394-0.500) 0.624 (0.562-0.674) 28.125 (9.343-50.876) 0.624 (0.441-0.751) 0.328 (0.217-0.425) 0.490 (0.340-0.603) 48.465 (12.673-92.137)
Gain $\uparrow$ +0.036 +0.012 +0.039 -39.748 +0.221 +0.103 +0.198 -80.448
Breast-only Unimodal 24 0.837 (0.787-0.878) 0.573 (0.516-0.627) 0.733 (0.668-0.790) 8.040 (4.885-11.685) 51 0.739 (0.705-0.771) 0.452 (0.423-0.481) 0.599 (0.559-0.639) 139.910 (114.241-167.642)
Multimodal 0.891 (0.870-0.908) 0.625 (0.580-0.669) 0.806 (0.775-0.833) 2.985 (2.365-3.616) 0.786 (0.749-0.820) 0.441 (0.411-0.473) 0.664 (0.625-0.700) 29.320 (20.006-40.850)
Gain $\uparrow$ +0.054 +0.052 +0.073 -5.055 +0.047 -0.011 +0.065 -110.59
@@
Table 7: Quantitative comparison results for our RO-LMM’s clinical report summarization and radiotherapy strategy suggestion performance on the publicly available dataset.
@@
R-1 $\uparrow$ (CI) R-2 $\uparrow$ (CI) R-L $\uparrow$ (CI)
@@
Radiotherapy Strategy Suggestion